OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)] can someone plz help me

OpenStudy (anonymous):

answer is 25

OpenStudy (anonymous):

how did u get 25

OpenStudy (anonymous):

first find g(-2) by substituting -2 in place of X

OpenStudy (anonymous):

i know that part butafter u get the first part which is three and the second part -13 wat do u do to those two numbers

OpenStudy (anonymous):

g(-2)=-13 then u will have f(-13) substitute it in place of x in f(x)

OpenStudy (anonymous):

If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)].this is another example

OpenStudy (anonymous):

let me try solving it tell me if i do it right

OpenStudy (anonymous):

answer is 1 check it !

OpenStudy (anonymous):

howd u get it show me all togethher

OpenStudy (anonymous):

If f(x) = 3x − 2 and g(x) = 2x − 1, find (f − g)(2).

OpenStudy (anonymous):

(f-g)(2)= f(2)- g(2) i hope u know how to find f(2) and g(2)

OpenStudy (anonymous):

nope tats wat im tryn to learn

OpenStudy (anonymous):

f(x) is a function in terms of x if u need to find f(2) thats mean u have to substitute the value of x as 2 in the given equation

OpenStudy (anonymous):

wats the answer

OpenStudy (anonymous):

answer is 1

OpenStudy (anonymous):

how bout this one If r(x) = −4x − 1 and t(x) = 2x − 1, find (r + t)(7)

OpenStudy (anonymous):

btw ure awesome u saved my life thank u

OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1).

OpenStudy (anonymous):

(f+g)(-1)=f(-1)+g(-1)

OpenStudy (anonymous):

if f(x) = x + 2 and g(x) = −3x − 1, find g[f(4)]

OpenStudy (anonymous):

ure good

OpenStudy (anonymous):

-19 is answer

OpenStudy (anonymous):

wait -19 is answer to which question

OpenStudy (anonymous):

If r(x) = −4x − 1 and t(x) = 2x − 1, find (r + t)(7) If r(x) = −4x − 1 and t(x) = 2x − 1, find (r + t)(7) If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1). If f(x) = x + 2 and g(x) = −3x − 1, find g[f(4)

OpenStudy (anonymous):

how bout thteese